0 0
Add a comment Improve this question Transcribed image text
Answer #1

12) Energy = power x time

power = p = VI = V2/R =0.266 W

Energy = 0.266 x 2 x 60 = 32 J

13) resistor are connected in series

hence Power P1 = V2/(R+65)

when the 65 ohms is replaced by 20 ohms

Power P2 = V2/(R+20)

but given P2 = 2P1

1/(R+20) = 2/(R+65)

R + 65 = 2R + 40

R = 25 Ohms

14) Since the current is 2 A

Lets calculate the drop across the 15 ohm resistor

V' = 15 x 2 = 30V Since this is higher than the applied voltage, 2 A current is not passing through 15 ohm resistor.Hence the resistors are connected parallely

Now effective resistance R'= 15R/(R+15) for parallel connection

Now voltage = iR'

20 = 30R/(R+15)

20R +300 = 30R

R = 30 ohms